Answer the following problems.


2.) Determine the sum pf the real zeros of P(x) = 3x4 + 14x3 + 10x2 - 14x + 3.

3.) Find μ so that x - 3 is a factor of  μx3 - 13x3 - μx - 3.








1
Expert's answer
2021-02-01T11:56:14-0500

2)

"x=-3"


"3(-3)^4+14(-3)^3+10(-3)^2-14(-3)+3=0"

"3x^4+14x^3+10x^2-14x+3="

"=3x^3(x+3)+5x^2(x+3)-5x(x+3)+x+3="

"=3(x+3)(x^3+\dfrac{5}{3}x^2-\dfrac{5}{3}x+\dfrac{1}{3})"

"=3(x+3)(x^3-\dfrac{1}{3}x^2+2x^2-\dfrac{2}{3}x-x+\dfrac{1}{3})"

"=3(x+3)(x-\dfrac{1}{3})(x^2+2x-1)"

"=3(x+3)(x-\dfrac{1}{3})(x+1-\sqrt{2})(x+1+\sqrt{2})"

"x_1+x_2+x_3+x_4=-3+\dfrac{1}{3}-2=-\dfrac{14}{3}"

Vieta's formulas


"x_1+x_2+x_3+x_4=-\dfrac{a_3}{a_4}=-\dfrac{14}{3}"


"x_1x_2+x_1x_3+x_1x_4+x_2x_3+x_2x_4+x_3x_4"


"=\dfrac{a_2}{a_4}=\dfrac{10}{3}"


"x_1x_2x_3+x_1x_2x_4+x_1x_3x_4+x_2x_3x_4"




"=-\dfrac{a_1}{a_4}=\dfrac{14}{3}"


"x_1x_2x_3x_4=\dfrac{a_0}{a_4}=1"

3)


"P(x)=\mu x^3-13x^2-\mu x-3"

"P(3)=\mu (3)^3-13(3)^2-\mu (3)-3=0"

"24\mu=120"

"\mu=5"